OpenStudy (anonymous):

\sqrt[3]{125x^14/y^3}

OpenStudy (saifoo.khan):

\[\sqrt[3]{\frac{125x^{14}}{y^3} }\]

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\Large \frac{5x^4 \sqrt[3]{y^2}}{y}\]

OpenStudy (anonymous):

( sorry tofu should i post all the probs cuz i gotta do my report on Africa)

OpenStudy (saifoo.khan):

let's do them one by one.

OpenStudy (anonymous):

O_o fine

OpenStudy (saifoo.khan):

do u have more?

OpenStudy (anonymous):

Yeah -.- I have Mathes mate

OpenStudy (anonymous):

evaluate the polynomial for x=5. when x=5, 3x^2-2x+4

OpenStudy (saifoo.khan):

69

OpenStudy (anonymous):

solve 25x^4-20x^2+3=0

OpenStudy (saifoo.khan):

not the whole final. :(

OpenStudy (anonymous):

no I did all but 15 of the problems

OpenStudy (saifoo.khan):

Ohh!

OpenStudy (anonymous):

I only need help with 2 more

OpenStudy (saifoo.khan):

alright. post them.

OpenStudy (saifoo.khan):

:)

OpenStudy (anonymous):

solve 25x^4-20x^2+3=0

OpenStudy (anonymous):

solve for x

OpenStudy (saifoo.khan):

 four answers.